Solve the equation w=Cr-2 for r where w is the weight of an object, C is a constant and r is the distance that the object is from the center of the earth.
Answer by stanbon(75887)

Solve the equation w=Cr-2 for r
-----------------
Cr = w+2
r = (w+2)/C
===============
Cheers,
Stan H.